ABOUT NATURE OF NONCOINCIDENCE FOR FREQUEHCY AND LINE WIDTH OF STOKES AND ANTISTOKES COMPONENTS IN RAMAN SPECTRA

\underline{E. N. SHERMATOV}, Samarkand State University, Univ. bulv 15, 703004 Samarkand, Uzbekistan.

Frequencies and widths (Δy_{1/2}) of isotropic and anisatropic (1) components of Raman lines was studied. In most cases are y>y and Δy_{1/2} > Δy_{1/2} (11). The study of stoces and antistoces components has been found the noncoincidence of line freguenc at given temperature, in most cases y_{i}^{ast} >y_{i}^{st} and Δy_{1/2}(ast)> Δy_{1/2} (st). Obtained results are explained in the context described by us author. Molecules in ground and excited vibration state are in different inner fields. There from natural frequecies of vibration in excited state is larger than those in ground state. This leads to relations y_{i}^{ast} >y_{i}^{st} and Δy_{1/2}(ast)> Δy_{1/2} (st).